### v2

# IndicGenBench: A Multilingual Benchmark to Evaluate Generation Capabilities of LLMs on Indic Languages

Abstract
As large language models (LLMs) see increasing adoption across the globe, it is imperative for LLMs to be representative of the linguistic diversity of the world. India is a linguistically diverse country of 1.4 Billion people. To facilitate research on multilingual LLM evaluation, we release IndicGenBench — the largest benchmark for evaluating LLMs on user-facing generation tasks across a diverse set 29 of Indic languages covering 13 scripts and 4 language families. IndicGenBench is composed of diverse generation tasks like cross-lingual summarization, machine translation, and cross-lingual question answering. IndicGenBench extends existing benchmarks to many Indic languages through human curation providing multi-way parallel evaluation data for many under-represented Indic languages for the first time. We evaluate a wide range of proprietary and open-source LLMs including GPT-3.5, GPT-4, PaLM-2, mT5, Gemma, BLOOM and LLaMA on IndicGenBench in a variety of settings. The largest PaLM-2 models performs the best on most tasks, however, there is a significant performance gap in all languages compared to English showing that further research is needed for the development of more inclusive multilingual language models. IndicGenBench is available at www.github.com/google-research-datasets/indic-gen-bench ††All authors are now part of Google DeepMind
IndicGenBench: A Multilingual Benchmark to Evaluate Generation Capabilities of LLMs on Indic Languages
Harman Singh, Nitish Gupta, Shikhar Bharadwaj, Dinesh Tewari, Partha Talukdar Google Research India {hrman, guptanitish, shikharop, dineshtewari, partha}@google.com
| Task | Language | Input | Output | #Languages | Dataset Size |
| (H / M / L) | (Train / Dev / Test) | ||||
| CrossSum-In (Cross-lingual Summarization) | Hindi |
9 / 7 / 13 | 2.9k / 2.9k / 14.5k |
| Flores-In (Machine Translation) | Konkani |
9 / 7 / 13 | - / 28.9k / 29.3k |
| XQuAD-In (Multilingual QA) | Punjabi |
9 / 3 / - | 1.2k / 1.2k / 14.3k |
| XorQA-In-Xx (Cross-lingual QA) | Telugu |
9 / 6 / 13 | 2.8k / 14k / 15.1k |
| XorQA-In-En (Cross-lingual QA) | Santali |
9 / 6 / 13 | 2.8k / 14k / 15.1k |
1 Introduction
With the advances in generative language technologies powered by Large Language Models (LLMs; Brown et al., 2020; Rae et al., 2021; Chowdhery et al., 2022; OpenAI et al., 2023; Tay et al., 2023; Google, 2023), there has been a surge of interest in evaluating the multilingual capabilities of these models. Recent work (Ahuja et al., 2023a, b) shows a consistent performance gap between high resource languages and languages with lower amounts of web resources available. To develop highly multilingual generative LLMs which should work equally well for 100s of languages spoken by billions of people in the world, it is crucial to evaluate their capabilities across a variety of languages to uncover performance gaps and guide future research.
In this work we focus on India, a country with 1369 rationalized mother tongues spoken by more than a billion people.111https://en.wikipedia.org/wiki/Languages_of_India Making progress on language technologies for Indic languages will not only improve the state of affairs in this region, but will also provide valuable learning to the NLP community which will be applicable to other geographical regions and language families. There are has been much work from the community in building natural language understanding (NLU) models for Indic languages (Kakwani et al., 2020; Khanuja et al., 2021), as well as evaluation datasets (Doddapaneni et al., 2023; Mhaske et al., 2023) to support such models. In this work, our focus is to develop a high-quality benchmark for evaluating generative language capabilities in a variety of Indic languages across various levels of resourcefulness.
We release IndicGenBench, a multilingual, multi-way parallel benchmark for measuring language generation capabilities across diverse user-facing tasks in 29 Indic languages across 4 language families (Table 7). IndicGenBench extends existing benchmarks such as CrossSum (Bhattacharjee et al., 2023), XQuAD (Artetxe et al., 2020), XorQA (Asai et al., 2021), and FLORES (NLLB-Team et al., 2022) for additional Indic languages and is composed of tasks like cross-lingual summarization (CrossSum-In), machine translation (Flores-In), cross-lingual reading comprehension (XorQA-In-Xx and XorQA-In-En) and multilingual reading comprehension (XQuAD-In). Each dataset consists of parallel examples in up to 29 low to comparatively higher resource Indic languages; and for some tasks (e.g. CrossSum-In), IndicGenBench provides the first-ever evaluation set for as many as 18 of these languages. We also release a small training set in all tasks for efficient adaptation of LLMs. Our comprehensive evaluation of various state-of-the-art proprietary and open-source LLMs on IndicGenBench shows that there is a significant gap in performance between English and Indic languages (see Figure 1). Our contributions are as follows:
Created and released IndicGenBench, a high quality text benchmark in diverse language generation tasks like summarization, question-answering, and translation across 29 Indic languages. IndicGenBench is the largest generation benchmark for Indic languages.
Comprehensive experimentation on SoTA LLMs (mT5, Gemma, BLOOM, LLaMA, GPT-3.5, GPT-4, PaLM-2) across various model sizes and training settings to benchmark their Indic language generation capabilities.
A qualitative analysis for assessing the gaps in current language technologies and define potential directions of future research.
2 IndicGenBench
IndicGenBench is a high-quality, human-curated benchmark to evaluate text generation capabilities of multilingual models on Indic languages. Our benchmark consists of 5 user-facing tasks (viz., summarization, machine translation, and question answering) across 29 Indic languages spanning 13 writing scripts and 4 language families. For certain tasks, IndicGenBench provides the first-ever evaluation dataset for up to 18 Indic languages. Table 1 provides summary of IndicGenBench and examples of instances across tasks present in it.
Languages in IndicGenBench are divided into (relatively) Higher, Medium, and Low resource categories based on the availability of web text resources (see appendix §A for details).222We note that the languages called relatively higher resource in this paper, e.g., Hindi or Bengali, are in fact mid-low Web resource when compared to English and other truly high resource languages. For example, using Wikipedia as a proxy for language resources, compared to 6.6M+ Wikipedia articles in English, there are only 160K Hindi Wikipedia articles.
Higher (9): Bengali, Gujarati, Hindi, Kannada, Malayalam, Marathi, Tamil, Telugu, Urdu
Medium (7): Assamese, Bhojpuri, Nepali, Odia, Punjabi, Pashto, Sanskrit
Low (13): Awadhi, Haryanvi, Tibetan, Garhwali, Konkani, Chhattisgarhi, Rajasthani, Maithili, Manipuri, Malvi, Marwari, Santali, Bodo
As evident from the lists above, our benchmark provides a broad-coverage over languages with respect to their resourcedness, allowing users to evaluate language models on relatively high-resource languages such as Hindi and extremely low-resource languages such as Manipuri in Meitei script on a single benchmark.
To curate the evaluation datasets for our benchmark, we use the following existing datasets as the source: CrossSum Bhattacharjee et al. (2023) for cross-lingual summarization, FLORES NLLB-Team et al. (2022) for machine translation, XQuAD Artetxe et al. (2020) for multilingual QA, and XoRQA Asai et al. (2021) for cross-lingual QA. From each of these datasets we select a subset of English examples to be a part of our benchmark, and then collect professional human translations for these examples in all target Indic languages. Some target languages are already covered by the source datasets in which case we re-purpose this existing data and only collect translations for the remaining languages. We also collect and release a small amount of training and validation data making possible evaluation of training techniques like fine-tuning, parameter-efficient training, in-context learning, and others.
Why extend existing benchmarks? We chose to collect human translations of existing benchmarks as opposed to creating evaluation data from scratch due to various reasons:
Translation-based extension of existing benchmark results in multi-way parallel data, allowing researchers to attribute performance due to task knowledge vs. language understanding, and measure cross-lingual generalization
For many low-resource languages in IndicGenBench, clean text knowledge corpus (e.g., Wikipedia) is not available making it difficult to acquire source data for annotation
By focusing only on translation quality in the target Indic languages, we are able to leverage the quality control that went into designing the source benchmarks.
Annotators were professional data labelers working as contractors at our organization and with a vendor. Annotators were paid competitive rates in compliance with applicable labor laws and prevailing market rates. Our pay rate to annotators varied across languages, ranging from USD 2.80 per hour for Pashto to USD 15.90 per hour for Tibetan.
Cross-Lingual Summarization: CrossSum-In
We create CrossSum-In based on CrossSum Bhattacharjee et al. (2023), a dataset for cross-lingual summarization, which in turn is derived from XL-Sum (Hasan et al., 2021b). CrossSum contains multi-way parallel data in 45 languages where BBC news articles as source in a language are paired with corresponding summaries in other languages. Based on their matching criteria, different languages have different amount of source-target pairs.
We sample 700 English article-summary pairs (100 each from train/dev and 500 from test) and ask human translators to translate the English summary into the target Indic languages. CrossSum already contains data for 9 of our 29 target languages; for these languages we sample 100/100/500 examples from the original dataset to maintain equity with other languages we collect data for. CrossSum-In contains a total of 20.3k examples across 29 Indic languages in our benchmark.
Machine Translation: Flores-In
FLORES-200 (NLLB-Team et al., 2022) is a human-annotated multi-way parallel machine translation (MT) benchmark for 200 languages where the same source English sentences are translated by humans into the target 200 languages. It contains data in 22 of our 29 target languages; we extend this by collecting human translations for the remaining 7 new languages leading to a MT benchmark in 29 Indic languages which we call Flores-In.
FLORES-200 is divided into three splits: dev (997), devtest (1012), test (992), of which the test set it not public. We collect translations for all 997 dev and 1012 devtest sentences, yielding 2009 sentences per language. Collectively, Flores-In contains 58.2k examples across 29 Indic languages.
Multilingual Question-Answering: XQuAD-In
We create an Indic Multilingual Question Answering task XQuAD-In based on the multilingual reading comprehension dataset XQuAD (Artetxe et al., 2020). XQuAD is in turn derived from the SQuAD dataset (Rajpurkar et al., 2016), in which an English Wikipedia passage is paired with multiple question-answer (QA) pairs where the answers are short spans for the given passage. The authors of XQuAD collected human translations for 240 passages and 1190 QA pairs from the SQuAD v1.1 development set into 10 higher resource languages (Hindi being the only Indic language).
To create XQuAD-In, we use the 240 passages and 1190 QA pairs from XQuAD as our test set. We additionally selected 20 passages and 100 QA pairs from the original SQuAD v1.1 training and development sets each to create our training and development set. For all the 280 passages and 1390 QA pairs we collect professional human translations in 12 Indic languages.333XQuAD-In contains all 9 higher-resource languages (see §2) and 3 medium-resources languages, namely, Assamese, Odia, and Punjabi. Overall, XQuAD-In contains 3.3k passages and 16.6k QA pairs in 12 Indic languages.
Cross-Lingual Question-Answering: XorQA-In
We create Indic Cross-lingual Question-Answering dataset XorQA-In based on the Xor-TyDi QA dataset (Asai et al., 2021). Xor-TyDi contains questions in non-English languages paired with English evidence passages and short span answers from those passages (similar to SQuAD). It was created with the idea of developing NLP systems that can answer questions in users’ native language by refering to sources in a high-resource language, such as English, which was more likely to contain the answer due to the information scarcity of low-resources languages on the web. The original Xor-TyDi contains data in 7 languages out of which Bengali and Telugu are the two Indic languages.
To create XorQA-In, we select the 302 Bengali and 237 Telugu examples (Bn/Te-question, En-passage, En-answer) from the Xor-TyDi dev set as our test data.444Xor-TyDi has not publicly released its test set. Additionally, we sample 600 examples (equally from Bengali and Telugu) from the training set of Xor-TyDi to create our training (100) and development (500) set. We then follow a two-staged translation process, where we first ask the human translators to translate the Bengali or Telugu question (Bn/Te-question) into English (En-question). In the second stage, we collect translations for these English questions (En-question) into target languages (Xx-question) and translations for the English answers (En-answer) into the target languages (Xx-answer).
We create two tasks from this translated data:
1. XorQA-In-En: Each example contains (Xx-question, En-passage, En-answer). This task is similar to the Xor-TyDi dataset.
2. XorQA-In-Xx: Each example contains (Xx-question, En-passage, Xx-answer), where the task is to generate the answer in the same language as the question.
We collect data for 28 Indic languages resulting in 32k examples.555We do not collect translations for Nepali.
See Appendix Table 9 for languages covered by each dataset in IndicGenBench.
| Model (LLM) |
CrossSum-In |
Flores-In |
XQuAD-In |
XorQA-In-Xx |
XorQA-In-En |
| Eval. Metric | ChrF | ChrF | Token-F1 | Token-F1 | Token-F1 |
| (enxx / xxen) | |||||
| Performance in English | |||||
| GPT-4 | 30.3 | – / – | 64.8 | – | 37.9 |
| PaLM-2-L | 41.1 | – / – | 83.7 | – | 71.4 |
| Average Performance on IndicGenBench | |||||
| LLaMA-7B | 3.7 | 11.5 / 21.6 | 3.8 | 7.4 | 10.4 |
| LLaMA-13B | 4.1 | 13.3 / 24.1 | 4.5 | 10.4 | 12.1 |
| LLaMA-65B | 4.6 | 18.1 / 32.7 | 7.1 | 16.5 | 16.3 |
| BLOOM-7B | 3.8 | 18.3 / 31.2 | 13.8 | 7.9 | 23.6 |
| BLOOMZ-7B | 1.2 | 40.8 / 48.4 | 53.7 | 7.0 | 49.0 |
| Gemma-7B-PT | 0.0 | 32.1 / 50.4 | 0.5 | 11.7 | 23.8 |
| Gemma-7B-IT | 11.6 | 18.6 / 29.2 | 35.3 | 13.5 | 24.8 |
| GPT-3.5 | 16.3 | 29.2 / 47.7 | 33.2 | 21.6 | 35.5 |
| GPT-4 | 17.6 | 32.1 / 54.5 | 55.7 | 23.4 | 46.0 |
| PaLM-2-XXS | 7.2 | 24.0 / 43.4 | 34.6 | 13.5 | 36.8 |
| PaLM-2-XS | 15.5 | 40.7 / 58.3 | 62.2 | 29.5 | 47.8 |
| PaLM-2-S | 18.5 | 43.5 / 61.6 | 66.7 | 31.6 | 57.4 |
| PaLM-2-L | 21.2 | 47.5 / 65.1 | 69.3 | 37.4 | 55.9 |
| CrossSum-In | Flores-In (enxx / xxen) | XQuAD-In | XorQA-In-Xx | XorQA-In-En | ||||||||||
| Model | High | Medium | Low | High | Medium | Low | High | Medium | High | Medium | Low | High | Medium | Low |
| LLaMA-65B | 4.4 | 4.6 | 4.7 | 18.2 / 31.5 | 15.4 / 30.0 | 19.5 / 35.0 | 8.8 | 1.9 | 17.7 | 13.5 | 17.1 | 16.4 | 14.0 | 17.3 |
| Gemma-7B-IT | 13.9 | 11.5 | 10.0 | 17.6 / 33.7 | 15.0 / 26.1 | 21.3 / 27.7 | 38.8 | 24.8 | 18.9 | 8.3 | 12.2 | 29.5 | 23.9 | 21.9 |
| BLOOMZ-7B | 1.5 | 1.7 | 0.6 | 67.7 / 59.1 | 39.4 / 50.2 | 22.9 / 40.0 | 55.5 | 48.1 | 10.8 | 2.8 | 6.2 | 64.7 | 45.8 | 39.5 |
| GPT-4 | 19.4 | 17.9 | 16.3 | 36.2 / 59.6 | 30.7 / 55.2 | 29.9 / 50.5 | 56.1 | 54.6 | 25.8 | 21.6 | 22.6 | 49.4 | 50.0 | 41.8 |
| PaLM-2-L | 25.2 | 23.1 | 17.5 | 56.9 / 68.2 | 45.9 / 65.6 | 41.9 / 62.6 | 72.5 | 59.8 | 41.9 | 36.7 | 34.6 | 57.3 | 57.9 | 53.9 |
3 Experiments and Analysis
We use IndicGenBench to benchmark multilingual and cross-lingual language generation capabilities of various LLMs on Indic languages. We perform experiments with a variety of open-source LLMs — mT5 (Xue et al., 2021), LLaMA (Touvron et al., 2023),666LLaMA-2 could not be used due to a restrictive licence, BLOOMZ (Workshop et al., 2022), Gemma (Team et al., 2024); and proprietary LLMs — GPT-3.5, GPT-4 (OpenAI et al., 2023), and PaLM-2 (Anil et al., 2023).
We compare and analyze the performance of different model size variants of these LLMs under various learning paradigm settings. We first evaluate model performance on one-shot prompting (§3.1) and also measure performance across language categories based on resourcedness (§3.2). We then evaluate the effect of number of in-context examples shown to the model as supervised data (§3.3) and the effect of prompting in a higher-resource language such an English or Hindi (§3.4). Using the training data contained in IndicGenBench, we measure how the performance of LLMs after fine-tuning compares with few-shot prompting (§3.5). Finally, we perform qualitative analysis of models on IndicGenBench and highlight some areas of improvement for future model development (§3.7).
Evaluation Metrics
For the cross-lingual summarization and translation tasks, CrossSum-In and Flores-In, we report Character-F1 (ChrF) metric (Popović, 2015) since token-level metrics like ROUGE and BLEU are not reliable for low-resource languages (Bapna et al., 2022). To stay consistent with existing literature on QA tasks, we report SQuAD-style Token-F1 on our XQuAD-In and XorQA-In QA tasks.
On Flores-In, we report translation performance in both directions—translating from English to the target language (enxx) and vice-versa (xxen).
3.1 Comparison of LLMs on IndicGenBench
In Table 2 we evaluate LLaMA, BLOOMZ, Gemma, GPT and PaLM-2 family of models on all tasks of IndicGenBench in a one-shot prompted setting. Numbers are averaged across all languages in the evaluation data. To compare, we also report English performance for GPT-4 and PaLM-2-L.
We see across tasks that larger models from the same LLM family perform better. PaLM-2-L performs the best among all LLMs considered, except for the XorQA-In-En task where PaLM-2-S performs slightly better. We find that open source LLaMA models perform much worse compared to proprietary models; even the largest LLaMA-65B model significantly underperforms the smallest PaLM-2-XXS model. Gemma-7B instruction tuned model performs better than LLaMA-13B as well as LLaMA-65B on most tasks. BLOOMZ, which is an instruction tuned version of BLOOM (Workshop et al., 2022), pre-trained on large-scale multilingual data, works the best on three out of five tasks in IndicGenBench. On CrossSum-In and XorQA-In-Xx it falls behind LLaMA and Gemma. Compared to English, we see significant room for improvement (20+ ChrF or Token-F1 points) across all tasks.
3.2 Performance across language categories
In Table 3 we report one-shot performance across language categories defined in Section 2. We only show performance for Gemma-7B-IT, BLOOMZ-7B, LLaMA-65B, GPT-4 and PaLM-2-L models here and report performance for the other models in appendix B.1. We find that there is a significant performance drop going from higher resourced languages to medium resourced ones, and further drop in lower resourced languages.
We would like to point out two observations here:
(a) In Flores-In, the performance for translating English to the target language (enxx) drops significantly from higher to lower resourced languages ( for PaLM-2-L) whereas the performance in the xxen direction does not fall this drastically (). This is also seen in XorQA-In-Xx and XorQA-In-En. This highlights that current LLMs are better at understanding than generation in these lower-resourced languages.
(b) In few cases, we see smaller performance deltas between medium and lower resourced languages compared to higher and medium categories. From our analysis, this can mainly be attributed to many languages in the lower category being similar to Hindi and written in the same Devanagari script.
| Flores-In | XorQA-In-Xx | ||||||
| Model (LLM) | 0 | 1 | 5 | 0 | 1 | 2 | 3 |
| LLaMA-7B | 8.0 | 11.5 | 11.4 | 5.0 | 7.4 | 9.0 | 9.2 |
| LLaMA-13B | 8.6 | 13.3 | 13.4 | 6.3 | 10.4 | 12.2 | 13.1 |
| LLaMA-65B | 14.0 | 18.1 | 18.3 | 12.3 | 16.5 | 18.7 | 19.4 |
| PaLM-2-XXS | 0.8 | 24.0 | 26.9 | 8.9 | 13.5 | 15.8 | 17.5 |
| PaLM-2-XS | 20.1 | 40.7 | 42.3 | 21.4 | 29.5 | 32.2 | 33.2 |
| PaLM-2-S | 24.9 | 43.5 | 45.2 | 22.7 | 31.6 | 33.4 | 35.4 |
| PaLM-2-L | 31.1 | 47.5 | 49.3 | 31.9 | 37.4 | 39.7 | 41.1 |
| CrossSum-In | XQuAD-In | XorQA-In-Xx | XorQA-In-En | ||||||||
| Model (1-Shot Lang) | Higher | Medium | Low | Higher | Medium | Higher | Medium | Low | Higher | Medium | Low |
| PaLM-2-XXS (En) | 0.3 | 0.1 | 0.3 | 38.5 | 31.9 | 14.0 | 5.4 | 7.3 | 40.3 | 35.0 | 30.8 |
| PaLM-2-XXS (Hi) | 1.3 | 2.1 | 3.7 | 39.8 | 33.3 | 17.6 | 8.5 | 10.5 | 45.5 | 39.4 | 31.9 |
| PaLM-2-XXS (Lang) | 7.7 | 7.6 | 6.7 | 37.2 | 26.8 | 17.7 | 8.8 | 12.8 | 43.6 | 38.3 | 31.5 |
| PaLM-2-XS (En) | 0.3 | 0.2 | 0.5 | 64.3 | 62.2 | 30.6 | 23.9 | 20.8 | 35.9 | 32.1 | 27.2 |
| PaLM-2-XS (Hi) | 3.5 | 5.5 | 9.9 | 65.4 | 63.5 | 33.2 | 25.8 | 22.7 | 49.3 | 46.8 | 40.7 |
| PaLM-2-XS (Lang) | 18.4 | 16.4 | 13.0 | 65.1 | 53.3 | 35.8 | 27.6 | 26.1 | 53.3 | 51.5 | 42.2 |
| PaLM-2-S (En) | 0.4 | 0.2 | 0.5 | 67.4 | 66.8 | 27.5 | 19.9 | 19.9 | 48.6 | 47.1 | 40.8 |
| PaLM-2-S (Hi) | 4.4 | 6.9 | 13.2 | 68.5 | 67.5 | 34.2 | 27.0 | 24.9 | 58.3 | 57.0 | 49.0 |
| PaLM-2-S (Lang) | 22.4 | 19.8 | 15.1 | 69.9 | 57.3 | 36.6 | 30.3 | 28.6 | 60.1 | 61.4 | 53.6 |
| PaLM-2-L (En) | 0.4 | 0.2 | 0.6 | 71.7 | 69.8 | 37.7 | 33.2 | 29.7 | 28.7 | 27.5 | 26.2 |
| PaLM-2-L (Hi) | 4.7 | 7.0 | 13.8 | 72.6 | 71.0 | 39.7 | 34.6 | 31.2 | 45.5 | 44.8 | 41.5 |
| PaLM-2-L (Lang) | 25.2 | 23.1 | 17.5 | 72.5 | 59.8 | 41.9 | 36.7 | 34.6 | 57.3 | 57.9 | 53.9 |
3.3 In-context learning on IndicGenBench
In this section we aim to understand the impact of the number of in-context examples shown to the LLM during few-shot prompting.
Since CrossSum-In and XQuAD-In input passages are long, we are only able to perform 0-and-1-shot prompting. For XorQA-In-Xx and XorQA-In-En we perform 0-to-3-shot prompting, and for Flores-In we perform 0, 1 and 5-shot prompting.
We show performance for Flores-In and XorQA-In-Xx in Table 4. Other results are shown in appendix 10 due to space limitations. Across model families and sizes we observe that increasing the amount of supervision in terms of the in-context examples improves performance.
| CrossSum-In | XQuAD-In | XorQA-In-Xx | XorQA-In-En | ||||||||
| Model | Higher | Medium | Low | Higher | Medium | Higher | Medium | Low | Higher | Medium | Low |
| mT5 models – Fine-Tuned | |||||||||||
| mT5-B | 19.5 | 18.9 | 15.1 | 46.2 | 30.9 | 3.8 | 4.0 | 5.5 | 31.7 | 31.4 | 30.8 |
| mT5-L | 20.5 | 19.9 | 15.5 | 54.3 | 38.6 | 11.8 | 11.0 | 10.4 | 56.8 | 53.7 | 45.4 |
| mT5-XL | 22.7 | 21.1 | 15.3 | 57.4 | 40.5 | 20.7 | 13.5 | 15.6 | 58.2 | 56.2 | 46.5 |
| mT5-XXL | 25.9 | 24.2 | 10.4 | 62.0 | 44.4 | 28.8 | 23.6 | 21.9 | 70.3 | 68.9 | 59.1 |
| PaLM-2 models - Fine-Tuned | |||||||||||
| PaLM-2-XXS | 22.5 | 19.7 | 16.5 | 41.2 | 18.1 | 18.1 | 10.9 | 12.9 | 60.2 | 56.9 | 50.9 |
| PaLM-2-XS | 28.5 | 25.6 | 18.8 | 40.2 | 16.9 | 30.4 | 23.6 | 19.6 | 69.1 | 66.6 | 56.6 |
| PaLM-2 models - Few-shot prompted | |||||||||||
| PaLM-2-XXSFS | 7.7 | 7.6 | 6.7 | 37.2 | 26.8 | 22.7 | 12.3 | 16.4 | 51.6 | 47.1 | 38.4 |
| PaLM-2-XSFS | 18.4 | 16.4 | 13.0 | 65.1 | 53.3 | 39.2 | 32.0 | 29.5 | 67.0 | 65.3 | 56.5 |
3.4 Transfer from high-resource languages
For languages with no supervised data, one option to improve performance is utilizing existing supervised data another language as in-context exemplars. In this section we aim to study if the language in which the model is prompted plays a role in performance.
In Table 5 we show performance when the model is prompted in English vs. Hindi, a representative higher resourced Indic language. For comparison, we also show performance when the in-context exemplar is in the same language as the test instance. We find that Hindi in-context exemplars are much more useful for all models as compared to their English counterparts. Surprisingly, for smaller models, performance with Hindi exemplars comes extremely close to prompting in the test language, even better sometimes.
3.5 Fine-tuning LLMs on IndicGenBench and Comparison with In-Context Learning
As outlined in Section 2, we also release a small, high-quality training set for all tasks in IndicGenBench (except Flores-In which only has dev and test sets). This training data can be used to adapt LLMs to downstream tasks in Indic languages via fine-tuning and other training techniques.
Table 6 shows our results of fine-tuning mT5 and PaLM-2 models and their comparison with in-context learning using PaLM-2. We fine-tune each model on training data from all available languages including English, use the development set for early stopping, and report numbers on the test set. For question-answering tasks that require generating short spans as answers, we find that older generation mT5 models significantly outperform smaller PaLM-2 models in most cases.777Since the parameter count for PaLM-2 models is not public, we cannot attribute this performance difference to model sizes. On CrossSum-In which requires generating a longer summary, we find that PaLM-2 models are more effective.
For Question-Answering tasks, as the model size increases from PaLM-2-XXS to PaLM-2-XS, we see that in-context learning yields equal or better performance compared to fine-tuning the model. For example, in XorQA-In-Xx, as the model size increases from XXS to XS, we see that the gap between few-shot prompting and fine-tuning significantly increases from 2-4% (in XXS) to 9-10% (in XS). In the case of XQuAD-In, we see that for the larger PaLM-2-XS model, its much better to perform in-context learning as compared to fine-tuning, for both medium and high resource Indic languages. For XorQA-In-En, in-context learning reaches the fine-tuning performance as model size increases to PaLM-2-XS. For the CrossSum-In, the gap between fine-tuning and in-context learning is reducing as model size increases, which reinforces that for even larger model sizes, it might be better to learn in-context.
3.6 Analyzing Tokenizer across Indic languages
In Figure 2, we compare the token fertility (average number of sub-words that a word is broken down into by the tokenizer) across all Indic langugaes in IndicGenBench.888We use OpenAI’s BPE tokenizer (platform.openai.com/tokenizer). PaLM-2 tokenizer is not publicly available. We find that the token fertility varies significantly across languages; from 4.1 for Pashto to 19.9 for Tibetan.
A high token fertility is undesirable and can disproportionately effect a particular language’s performance. For languages where text is broken into more number of tokens, fewer in-context examples can be input to the LLM during inference. This can negatively impact performance (see Table 4). In Figure 3, we show how the percentage of a dataset that fits in a particular context length changes with number of in-context examples for various languages. For example, we see in Figure 3 that for medium resource languages with high token-fertility like Oriya and Punjabi we can in-corporate much fewer in-context examples compared to Indic languages with lower token-fertility like Hindi and Marathi. Analysis in this section is inspired by prior work on studying the impact of tokenization on model utility, inference, and financial cost (Ahia et al., 2023; Petrov et al., 2023). Our work supports and compliments these studies for a range of Indic languages. Alternate tokenization and encoding approaches are required to bridge the tokenization disparity between different languages. An example of this is MYTE (Limisiewicz et al., 2024), a morphology based byte encoding scheme which leads to more equitable text representations.
3.7 Qualitative Analysis
We manually analyze predictions from the best performing model PaLM-2-L with the aim to understand the shortcomings of current LLMs and highlight areas of improvements for future research. We randomly select 20 examples each in the CrossSum-In and Flores-In tasks for the following languages which are reviewed by native speakers: Awadhi, Haryanvi, Chhatisgarhi, Konkani, and Assamese. We found the following patterns of errors:
Generation in a related language
The languages Awadhi, Haryanvi, and Chhatisgarhi are related to a higher resource language Hindi and written in the same script Devanagari. We find that the model generates mixed-language output with words mixed from Hindi and also outputs incorrectly inflected forms of the main verbs in the output. We show couple of examples of this phenomena in Figure 5(a) in the appendix.
Hallucination and Missing Information
In the cross-lingual summarization task CrossSum-In, we find that the model often outputs extra information that is not present in the source article. In translation, we have observed examples where come crucial information from the source sentence is missing from the generated output. Also, in some cases, the model fails to understand polysemous English words and generates translation for the incorrect sense. We show examples of these phenomena in Figures 4(a), 4(b), and 5(b) in the appendix.
4 Related Work
In the last few years, many multilingual LLMs have been developed—starting from mBART (Liu et al., 2020) trained on 25 languages to LLMs that are pre-trained on hundreds of languages, such as mT5 (Xue et al., 2021), PaLM-2 (Anil et al., 2023), GPT-4 (Achiam et al., 2023), Gemini (Google, 2023), and others. These LLMs are typically evaluated on individual multilingual tasks for Translation: WMT (Farhad et al., 2021), FLORES (NLLB-Team et al., 2022); Question-Answering: XQuAD (Artetxe et al., 2020), TyDiQA (Clark et al., 2020), XorQA (Asai et al., 2021); Summarization: XLSUM (Hasan et al., 2021a); Reasoning: MGSM (Shi et al., 2022), XCOPA (Ponti et al., 2020) to name a few, or on multilingual benchmarks such as, XTREME (Hu et al., 2020) and XTREME-UP (Ruder et al., 2023). However, most of these evaluation resources contain only a handful of languages or do not contain data for low resource languages, especially Indics. Besides, cross-lingual evaluation data is even more sparse. This work is an effort to bridge these gaps by releasing IndicGenBench, a suite of datasets covering diverse cross-lingual and multilingual generation tasks in Indic languages.
Most work on creating evaluation data on Indic languages have focused on natural language understanding (NLU) tasks. Kakwani et al. (2020) and Doddapaneni et al. (2023) have released NLU test sets in Indic languages for a wide variety of tasks such as QA and NLI. Naamapadam (Mhaske et al., 2023) is a named entity recognition dataset specifically for Indic languages, MASSIVE (FitzGerald et al., 2022) is a slot-filling and intent classification dataset available in 7 Indic languages, IndicGLUE (Kakwani et al., 2020) is an NLU benchmark for 11 Indic languages, whereas GLUECoS (Khanuja et al., 2020) is a Hindi-English code-mixed benchmark, containing various NLU tasks. The Belebele Benchmark (Bandarkar et al., 2023) is a multiple-choice machine reading comprehension dataset for 122 languages of which 17 are Indic. On the other hand, IndicGenBench is a natural language generation (NLG) benchmark.
Recently, there has been work in creating evaluation benchmarks for natural language generation (NLG) on Indic languages. IndicNLG Suite (Kumar et al., 2022), consisting of 5 NLG taks in 11 Indic languages, is a leap in this direction. These datasets in this suite are automatically created, either using data from the web (e.g., Wikipedia) or using translation systems. There are few works which create evaluation data for individual tasks in Indic languages. For example, IndicTrans2 (Gala et al., 2023) creates an n-way parallel dataset for machine translation in 22 scheduled Indian Languages, Mukhyansh (Madasu et al., 2023) and PMIndiaSum (Urlana et al., 2023) are headline generation datasets for 8 and 14 Indic languages respectively, and TeSum (Urlana et al., 2022) is an abstractive summarization dataset in the Telugu language. Ramesh et al. (2022) introduced Samanantar, a large translation dataset covering 11 Indic languages. Our work complements IndicNLGSuite and the other datasets in multiple ways. IndicGenBench is manually annotated ensuring high-quality, noise-free text which is not typically found on the web. Our benchmark contains evaluation data for a much larger set of languages spanning low, medium and high resource. Our datasets are multi-language parallel enabling better comparison among different languages. Lastly, we focus on a complementary and challenging set of tasks, including cross-lingual summarization, cross-lingual and multilingual question answering, and translation.
5 Conclusion
We release IndicGenBench, the largest benchmark for evaluating LLMs on 5 user-facing generation tasks across 29 Indic languages, providing evaluation data for many under-represented Indic languages for the first time. IndicGenBench is broad coverage along many dimensions – it covers 13 writing scripts, 4 language families, and spans languages across the available web resource spectrum. We carry out extensive comparison of current SoTA LLMs on IndicGenBench and highlight areas for future improvement. We are hopeful IndicGenBench will play an important role in further development of LLMs in Indic languages ultimately benefiting a billion-plus population.
6 Limitations
Since IndicGenBench extends existing benchmarks to new Indic languages through human translation, it may miss some India-specific entities and linguistic nuances. Future work can explore trans-localization for creating improved evaluation and fine-tuning. IndicGenBench doesn’t cover long-form generation and reasoning tasks. Creating such datasets is part of our future work.
Acknowledgments
We thank Aditi Chaudhury, Ashok Popat, Shachi Dave, Sagar Gubbi, Megh Umekar and members of the Languages team at Google Research India (GRI) for providing feedback on this work. The authors would like to thank Manish Gupta and Divy Thakkar for their support and guidance.
Appendix
| Code | Language | Script | Family | Resource |
| en | English | Latin | Germanic | high |
| bn | Bengali | Bengali | Indo-European | higher (Indic relative) |
| gu | Gujarati | Gujarati | Indo-European | higher (Indic relative) |
| hi | Hindi | Devanagari | Indo-European | higher (Indic relative) |
| kn | Kannada | Kannada | Dravidian | higher (Indic relative) |
| ml | Malayalam | Malayalam | Dravidian | higher (Indic relative) |
| mr | Marathi | Devanagari | Indo-European | higher (Indic relative) |
| ta | Tamil | Tamil | Dravidian | higher (Indic relative) |
| te | Telugu | Telugu | Dravidian | higher (Indic relative) |
| ur | Urdu | Arabic | Indo-European | higher (Indic relative) |
| as | Assamese | Bengali | Indo-European | med |
| bho | Bhojpuri | Devanagari | Indo-European | med |
| ne | Nepali | Devanagari | Indo-European | med |
| or | Odia | Odia | Indo-European | med |
| pa | Punjabi | Gurumukhi | Indo-European | med |
| ps | Pashto | Arabic | Indo-European | med |
| sa | Sanskrit | Devanagari | Indo-European | med |
| awa | Awadhi | Devanagari | Indo-European | low |
| bgc | Haryanvi | Devanagari | Indo-European | low |
| bo | Tibetan | Tibetan | Sino-Tibetan | low |
| brx | Bodo | Devanagari | Sino-Tibetan | low |
| gbm | Garhwali | Devanagari | Indo-European | low |
| gom | Konkani | Devanagari | Indo-European | low |
| hne | Chhattisgarhi | Devanagari | Indo-European | low |
| hoj | Rajasthani | Devanagari | Indo-European | low |
| mai | Maithili | Devanagari | Indo-European | low |
| mni | Manipuri | Meithi | Sino-Tibetan | low |
| mup | Malvi | Devanagari | Indo-European | low |
| mwr | Marwari | Devanagari | Indo-European | low |
| sat | Santali | Ol Chiki | Austroasiatic | low |
Appendix A Resource-wise Language Classification
See Table 7 for the language resource based classification of the languages studied in our work. This classification is done on the basis of multiple criteria, such as whether google translate supports the particular language, the classification of Indic languages suggested by Doddapaneni et al. (2023) which in turn uses Joshi et al. (2020).
Appendix B Detailed and Additonal Experimental Results
In this section, we report (a) one-shot results across all models and tasks averaged over different language categories (B.1); (b) performance for different tasks and models by varying the number of in-context exemplars shown in few-shot prompting (B.2); (c) one-shot performance for the largest model in each LLM family across all languages (B.3); (d) fine-tuning performance for the largest models across all languages (B.4).
B.1 Resource wise one-shot results
In Table 10 we show performance for all LLMs considered in this paper on the one-shot prompting method across different language categories.
B.2 Performance by varying number of in-context examples
B.3 Language wise one-shot performance across models
B.4 Language wise fine-tuning performance across models
B.5 Indic Specific LLMs
To the best of our knowledge, there is no LLM that is pre-trained or fine-tuned on a broad set of Indic languages, or a majority of the 29 languages we study in this work. Some Indic LLMs such as OpenHathi (or its instruction-tuned variant Airavata (Gala et al., 2024)), Tamil-LLama (Balachandran, 2023), or Kan-LLama are primarily trained/fine-tuned on a single Indic language (Hindi, Telugu, and Kannada, respectively).
In this section we evaluate an Indic specific LLM, Airavata which is primarily pre-trained and fine-tuned on English and Hindi language data. We evaluate the model on 6 Indic languages: Hindi (hi), Bengali (bn), Punjabi (pa), Assamese (as), Manipuri (mni) and Santali (sat) (2 each from higher, medium and low resource), and English. See Table 8 for the results.
We find that Airavata performs reasonably well only on the Hindi language, and its performance on other Indic languages is much worse (as expected since it is only trained for Hindi). Airavata performs significantly better than LLaMA on the Hindi language, across tasks, except for CSSUM-IN summarization tasks. BLOOMZ outperforms Airavata on most tasks in the Hindi language.
| CSSUM-IN | FLORES-IN (enxx) | XQUAD-IN | XORQA-IN-XX | XORQA-IN-EN | |||||||||||
| Language | Airavata | LLaMA | BLOOMZ | Airavata | LLaMA | BLOOMZ | Airavata | LLaMA | BLOOMZ | Airavata | LLaMA | BLOOMZ | Airavata | LLaMA | BLOOMZ |
| en | 31.6 | 19.1 | 13.9 | - | - | - | 75.7 | 49.2 | 86.8 | 69.5 | 56.0 | 72.2 | 69.5 | 59.7 | 68.7 |
| hi | 0.3 | 6.0 | 1.1 | 55.9 | 19.0 | 66.8 | 48.7 | 16.6 | 64.3 | 46.8 | 30.3 | 18.6 | 31.6 | 19.4 | 67.6 |
| bn | 0.0 | 2.0 | 0.2 | 8.5 | 9.6 | 73.9 | 6.5 | 2.5 | 57.9 | 0.2 | 3.8 | 0.3 | 33.8 | 12.6 | 70.0 |
| pa | 0.1 | 1.2 | 0.3 | 6.8 | 6.8 | 55.7 | 3.2 | 0.8 | 60.2 | 2.5 | 0.9 | 4.7 | 21.2 | 1.8 | 67.3 |
| as | 0.3 | 2.0 | 1.4 | 5.8 | 5.5 | 50.1 | 5.2 | 2.1 | 52.8 | 0.5 | 3.0 | 2.6 | 32.4 | 9.4 | 57.7 |
| mni | 0.2 | 2.3 | 0.2 | 6.1 | 6.7 | 6.5 | - | - | - | 1.4 | 2.9 | 0.3 | 22.6 | 5.4 | 21.5 |
| sat | 0.0 | 2.4 | 0.3 | 6.4 | 8.3 | 15.9 | - | - | - | 1.1 | 1.1 | 0.0 | 18.6 | 5.7 | 5.5 |
| Code | Language | Coverage in IndicGenBench | ||||
| CrossSum-In | Flores-In | XQuAD-In | XorQA-In-Xx | XorQA-In-En | ||
| bn | Bengali | ✓ | ✓ | ✓ | ✓ | ✓ |
| gu | Gujarati | ✓ | ✓ | ✓ | ✓ | ✓ |
| hi | Hindi | ✓ | ✓ | ✓ | ✓ | ✓ |
| kn | Kannada | ✓ | ✓ | ✓ | ✓ | ✓ |
| ml | Malayalam | ✓ | ✓ | ✓ | ✓ | ✓ |
| mr | Marathi | ✓ | ✓ | ✓ | ✓ | ✓ |
| ta | Tamil | ✓ | ✓ | ✓ | ✓ | ✓ |
| te | Telugu | ✓ | ✓ | ✓ | ✓ | ✓ |
| ur | Urdu | ✓ | ✓ | ✓ | ✓ | ✓ |
| as | Assamese | ✓ | ✓ | ✓ | ✓ | ✓ |
| bho | Bhojpuri | ✓ | ✓ | ✓ | ✓ | |
| ne | Nepali | ✓ | ✓ | |||
| or | Odia | ✓ | ✓ | ✓ | ✓ | ✓ |
| pa | Punjabi | ✓ | ✓ | ✓ | ✓ | ✓ |
| ps | Pashto | ✓ | ✓ | ✓ | ✓ | |
| sa | Sanskrit | ✓ | ✓ | ✓ | ✓ | |
| awa | Awadhi | ✓ | ✓ | ✓ | ✓ | |
| bgc | Haryanvi | ✓ | ✓ | ✓ | ✓ | |
| bo | Tibetan | ✓ | ✓ | ✓ | ✓ | |
| brx | Bodo | ✓ | ✓ | ✓ | ✓ | |
| gbm | Garhwali | ✓ | ✓ | ✓ | ✓ | |
| gom | Konkani | ✓ | ✓ | ✓ | ✓ | |
| hne | Chhattisgarhi | ✓ | ✓ | ✓ | ✓ | |
| hoj | Rajasthani | ✓ | ✓ | ✓ | ✓ | |
| mai | Maithili | ✓ | ✓ | ✓ | ✓ | |
| mni | Manipuri | ✓ | ✓ | ✓ | ✓ | |
| mup | Malvi | ✓ | ✓ | ✓ | ✓ | |
| mwr | Marwari | ✓ | ✓ | ✓ | ✓ | |
| sat | Santali | ✓ | ✓ | ✓ | ✓ | |
| XorQA-In-En | XQuAD-In | CrossSum-In | XorQA-In-Xx | Flores-In | ||||||||||
| Model | High | Medium | Low | High | Medium | High | Medium | Low | High | Medium | Low | High | Medium | Low |
| LLaMA-7B | 9.9 | 7.6 | 12.2 | 4.7 | 1.1 | 3.6 | 3.6 | 3.7 | 8.1 | 3.4 | 8.8 | 11.2/20.4 | 9.6/20.2 | 12.8/23.2 |
| LLaMA-13B | 10.5 | 9.9 | 14.2 | 5.5 | 1.5 | 3.9 | 4.1 | 4.2 | 9.6 | 6.3 | 12.8 | 12.6/22.5 | 11.3/22.0 | 14.8/26.4 |
| LLaMA-65B | 16.4 | 14.0 | 17.3 | 8.8 | 1.9 | 4.4 | 4.6 | 4.7 | 17.7 | 13.5 | 17.1 | 18.2/31.5 | 15.4/30.0 | 19.5/35.0 |
| Gemma-7B-PT | 25.7 | 22.5 | 23.1 | 0.6 | 0.1 | 0.0 | 0.0 | 0.0 | 16.5 | 6.9 | 10.5 | 40.7 / 58.8 | 25.5 / 49.5 | 29.8 / 45.0 |
| Gemma-7B-IT | 29.5 | 23.9 | 21.9 | 38.8 | 24.8 | 13.9 | 11.5 | 10.0 | 18.9 | 8.3 | 12.2 | 17.6 / 33.7 | 15.0 / 26.1 | 21.3 / 27.7 |
| BLOOM-7B | 29.3 | 21.6 | 20.6 | 15.2 | 9.8 | 3.1 | 3.8 | 4.3 | 10.8 | 5.0 | 7.2 | 19.8 / 34.7 | 16.3 / 31.4 | 18.4 / 28.6 |
| BLOOMZ-7B | 64.7 | 45.8 | 39.5 | 55.5 | 48.1 | 1.5 | 1.7 | 0.6 | 10.8 | 2.8 | 6.2 | 67.7 / 59.1 | 39.4 / 50.2 | 22.9 / 40.0 |
| GPT-3.5 | 39.0 | 34.9 | 33.4 | 36.2 | 23.9 | 17.6 | 16.5 | 15.3 | 24.2 | 20.0 | 20.6 | 32.9/52.9 | 27.3/48.2 | 27.6/43.8 |
| GPT-4 | 49.4 | 50.0 | 41.8 | 56.1 | 54.6 | 19.4 | 17.9 | 16.3 | 25.8 | 21.6 | 22.6 | 36.2/59.6 | 30.7/55.2 | 29.9/50.5 |
| PaLM-2-XXS | 43.6 | 38.3 | 31.5 | 37.2 | 26.8 | 7.7 | 7.6 | 6.7 | 17.7 | 8.8 | 12.8 | 31.7 / 52.1 | 19.0 / 43.2 | 21.4 / 37.6 |
| PaLM-2-XS | 53.3 | 51.5 | 42.2 | 65.1 | 53.3 | 18.4 | 16.4 | 13.0 | 35.8 | 27.6 | 26.1 | 52.0 / 64.8 | 39.0 / 60.6 | 33.7 / 52.6 |
| PaLM-2-S | 60.1 | 61.4 | 53.6 | 69.9 | 57.3 | 22.4 | 19.8 | 15.1 | 36.6 | 30.3 | 28.6 | 54.7 / 66.8 | 42.5 / 63.5 | 36.4 / 57.0 |
| PaLM-2-L | 57.3 | 57.9 | 53.9 | 72.5 | 59.8 | 25.2 | 23.1 | 17.5 | 41.9 | 36.7 | 34.6 | 56.9 / 68.2 | 45.9 / 65.6 | 41.9 / 62.6 |
| Shots | ||
| Model (LLM) | 0 | 1 |
| LLaMA-7B | 2.4 | 3.7 |
| LLaMA-13B | 1.5 | 4.1 |
| LLaMA-65B | 5.1 | 4.6 |
| PaLM-2-XXS | 0.2 | 7.2 |
| PaLM-2-XS | 1.0 | 15.5 |
| PaLM-2-S | 4.1 | 18.5 |
| PaLM-2-L | 7.7 | 21.2 |
| Shots | ||
| Model (LLM) | 0 | 1 |
| LLaMA-7B | 4.3 | 3.8 |
| LLaMA-13B | 4.7 | 4.5 |
| LLaMA-65B | 7.2 | 7.1 |
| PaLM-2-XXS | 19.0 | 34.6 |
| PaLM-2-XS | 38.7 | 62.2 |
| PaLM-2-S | 44.7 | 66.7 |
| PaLM-2-L | 50.6 | 69.3 |
| Shots | ||||
| Model (LLM) | 0 | 1 | 2 | 3 |
| LLaMA-7B | 4.7 | 10.4 | 12.5 | 14.1 |
| LLaMA-13B | 2.7 | 12.1 | 15.4 | 17.0 |
| LLaMA-65B | 8.2 | 16.3 | 18.8 | 19.8 |
| PaLM-2-XXS | 31.0 | 36.8 | 41.6 | 44.5 |
| PaLM-2-XS | 15.6 | 47.8 | 57.8 | 61.7 |
| PaLM-2-S | 40.6 | 57.4 | 63.4 | 66.3 |
| PaLM-2-L | 31.4 | 55.9 | 66.3 | 70.3 |
| # of in-context examples | |||
| Model (LLM) | 0 | 1 | 5 |
| LLaMA-7B | 8.0 / 10.4 | 11.5 / 21.6 | 11.4 / 19.8 |
| LLaMA-13B | 8.6 / 20.6 | 13.3 / 24.1 | 13.4 / 22.7 |
| LLaMA-65B | 14.0 / 28.0 | 18.1 / 32.7 | 18.3 / 30.3 |
| PaLM-2-XXS | 0.8 / 22.4 | 24.0 / 43.4 | 26.9 / 44.8 |
| PaLM-2-XS | 20.1 / 53.3 | 40.7 / 58.3 | 42.3 / 58.8 |
| PaLM-2-S | 24.9 / 60.1 | 43.5 / 61.6 | 45.2 / 61.9 |
| PaLM-2-L | 31.1 / 62.2 | 47.5 / 65.1 | 49.3 / 65.7 |
Appendix C Hyperparameters
For Fine-tuning experiments on mT5, we hyperparameter search for batch size in the range {16, 32, 64} and learning rate in {5e-3, 1e-3, 5e-4, 1e-4, 5e-5, 1e-5} and select the best hyperparameters based on the Token-F1 (for Question answering datasets) score or ChrF (for CrossSum-In, Flores-In experiments) score on the validation set of the dataset. For PaLM-2 fine-tuning experiments, we hyperparameter search for batch size in the range {16, 32, 64} and learning rate in {5e-4, 1e-4, 5e-5, 1e-5} and select the best hyperparameters based on the Token-F1 score or ChrF score, as in the case of mT5 experiments. We keep temperature=0 as the default setting while decoding using PaLM-2 and GPT family of models. For LLaMA models we increased the temperature to 0.1, since otherwise, the performance of LLaMA were coming out to be close to 0 for many of our datasets. For each dataset, we fix a prompt for all models since searching for a prompt would have led to a much larger compute cost for training and evaluation. We also choose in-context exemplars randomly from the training set (from the dev set in the case of Flores-In). This decision is a limitation of this work, since it has been shown that prompt and in-context exemplars can significantly impact an LLMs performance (Zhao et al., 2021b, a). The prompts we use are provided in §F. Searching for better prompts and exemplars is left as future work. For all fine-tuning runs, we just perform one run (as opposed to taking average of multiple runs) due to prohibitive costs of fine-tuning LLMs multiple times.
Appendix D Computing Infra
For mT5 fine-tuning runs, we use a combination of 4 to 64 TPU-V3, and for PaLM-2 we use up to 256 TPU-V4. All experiments take about 6 hours to 1-2 days of time depending upon the resources and size of the model being trained. We use NVIDIA-V100 16GB GPUs for evaluating open-source models like BLOOM, BLOOMZ, Airavata on IndicGenBench.
Appendix E Dataset Licenses
In compliance with licenses of the original datasets, we release our evaluation datasets under the following licenses: (a) XQuAD-In and Flores-In under the CC BY-SA 4.0 license; (b) CrossSum-In under the CC BY-NC-SA 4.0 license; and (d) XorQA-In under the MIT license.
Appendix F Prompts
The set of prompts used for evaluations on all datasets are shown in Table 15.
| Dataset | Prompt |
| CrossSum-In |
I will first show a news article in English and then provide a summary of it in the [Target Language Name] language.
Summarize the following article: [Article] Summary: |
| Flores-In, xxen |
Translate the following:
To English: [Sentence in Target Language] Output: |
| Flores-In, enxx |
Translate the following:
To [Target Language Name]: [Sentence in English] Output: |
| XQuAD-In |
[Passage in Target Language]
Q: [Question in Target Language] A: |
| XorQA-In-Xx |
Generate an answer in [Target Language Name] for the question based on the given passage:
[Passage in English] Q: [Question in Target Language] A: |
| XorQA-In-Xx |
Generate an answer in English for the question based on the given passage:
[Passage in English] Q: [Question in Target Language] A: |
| Lang. Code | CrossSum-In | ||||
| LLaMA-65B | Gemma-7B-IT | BLOOMZ-7B | GPT-4 | PaLM-2-L | |
| en | 26.1 | 24.8 | 18.6 | 30.3 | 41.1 |
| bn | 2.7 | 13.6 | 0.4 | 20.9 | 23.4 |
| gu | 2.1 | 10.3 | 1.0 | 16.5 | 19.2 |
| hi | 7.8 | 16.8 | 0.9 | 23.1 | 30.4 |
| kn | 2.9 | 13.6 | 0.9 | 15.1 | 25.4 |
| ml | 4.3 | 11.4 | 0.5 | 14.4 | 23.3 |
| mr | 4.1 | 13.6 | 0.6 | 22.4 | 25.4 |
| ta | 5.4 | 20.7 | 2.1 | 20.4 | 28.9 |
| te | 2.5 | 13.6 | 0.7 | 16.6 | 21.9 |
| ur | 7.5 | 13.0 | 6.4 | 24.9 | 28.9 |
| as | 2.7 | 12.5 | 1.8 | 17.3 | 24.3 |
| bho | 5.5 | 13.2 | 0.9 | 19.3 | 20.2 |
| ne | 7.4 | 16.8 | 2.0 | 23.2 | 29.1 |
| or | 2.5 | 4.3 | 0.3 | 10.7 | 23.4 |
| pa | 1.7 | 10.7 | 0.2 | 15.8 | 19.1 |
| ps | 7.0 | 9.2 | 6.3 | 19.9 | 25.2 |
| sa | 5.3 | 13.5 | 0.5 | 19.4 | 20.1 |
| awa | 5.1 | 12.8 | 0.6 | 18.2 | 19.4 |
| bgc | 5.3 | 12.5 | 0.7 | 17.9 | 20.2 |
| bo | 1.5 | 2.8 | 0.1 | 8.5 | 10.0 |
| brx | 4.9 | 2.8 | 0.4 | 15.5 | 11.0 |
| gbm | 5.8 | 11.4 | 2.0 | 16.2 | 16.2 |
| gom | 5.8 | 12.6 | 0.8 | 20.3 | 23.1 |
| hne | 5.5 | 13.4 | 0.9 | 19.2 | 22.0 |
| hoj | 5.2 | 12.5 | 0.3 | 17.7 | 17.7 |
| mai | 4.9 | 12.6 | 0.5 | 18.2 | 21.1 |
| mni | 3.3 | 4.4 | 0.2 | 13.2 | 15.6 |
| mup | 5.3 | 13.1 | 0.6 | 20.0 | 23.8 |
| mwr | 5.1 | 13.2 | 0.8 | 18.4 | 20.2 |
| sat | 2.7 | 5.5 | 0.2 | 8.5 | 7.4 |
| Avg. | 4.6 | 11.6 | 1.2 | 17.6 | 21.2 |
| High Avg. | 4.4 | 13.9 | 1.5 | 19.4 | 25.2 |
| Medium Avg. | 4.6 | 11.5 | 1.7 | 17.9 | 23.1 |
| Low Avg. | 4.7 | 10.0 | 0.6 | 16.3 | 17.5 |
| Lang. Code | Flores-In (enxx) | Flores-In (xxen) | ||||||||
| LLaMA-65B | Gemma-7B-IT | BLOOMZ-7B | GPT-4 | PaLM-2-L | LLaMA-65B | Gemma-7B-IT | BLOOMZ-7B | GPT-4 | PaLM-2-L | |
| bn | 18.7 | 27.3 | 71.9 | 40.0 | 54.0 | 37.4 | 35.1 | 60.4 | 59.9 | 66.8 |
| gu | 11.3 | 3.3 | 68.1 | 31.3 | 56.1 | 22.1 | 29.1 | 60.9 | 61.2 | 70.9 |
| hi | 33.2 | 35.7 | 65.2 | 48.6 | 60.2 | 53.7 | 48.4 | 57.3 | 65.2 | 70.5 |
| kn | 11.1 | 15.6 | 66.8 | 29.9 | 58.0 | 20.6 | 27.6 | 59.1 | 57.6 | 65.7 |
| ml | 14.5 | 7.1 | 66.2 | 28.4 | 59.7 | 22.8 | 28.8 | 60.6 | 58.6 | 68.0 |
| mr | 22.7 | 22.3 | 68.6 | 39.2 | 53.1 | 38.2 | 31.7 | 58.8 | 59.8 | 68.5 |
| ta | 16.8 | 29.8 | 75.2 | 34.5 | 60.0 | 24.5 | 30.5 | 55.7 | 53.9 | 65.7 |
| te | 10.7 | 16.1 | 68.8 | 30.7 | 60.5 | 21.7 | 33.0 | 56.1 | 58.1 | 70.5 |
| ur | 24.6 | 1.3 | 58.5 | 43.5 | 50.6 | 42.8 | 38.8 | 62.8 | 62.0 | 67.2 |
| awa | 26.4 | 29.1 | 32.1 | 38.8 | 49.9 | 46.1 | 36.9 | 56.3 | 61.8 | 69.3 |
| bgc | 26.6 | 28.9 | 31.1 | 38.0 | 49.8 | 45.1 | 33.3 | 55.7 | 63.5 | 72.1 |
| bo | 6.6 | 4.5 | 13.7 | 17.0 | 40.9 | 16.9 | 17.0 | 16.1 | 27.7 | 50.1 |
| brx | 12.4 | 13.0 | 12.1 | 17.2 | 11.3 | 20.7 | 17.8 | 17.0 | 29.2 | 31.0 |
| gbm | 24.0 | 26.6 | 28.3 | 36.2 | 47.9 | 43.0 | 31.2 | 49.6 | 62.2 | 72.1 |
| gom | 17.8 | 19.0 | 17.3 | 27.5 | 39.0 | 29.7 | 23.3 | 28.3 | 48.2 | 63.2 |
| hne | 25.4 | 29.2 | 29.3 | 38.4 | 52.7 | 42.8 | 32.9 | 53.6 | 62.3 | 75.1 |
| hoj | 23.7 | 26.6 | 28.1 | 36.0 | 48.1 | 41.9 | 30.4 | 51.3 | 62.3 | 73.0 |
| mai | 21.4 | 26.0 | 22.8 | 36.2 | 54.1 | 41.8 | 32.7 | 57.5 | 60.9 | 73.3 |
| mni | 8.3 | 9.3 | 8.7 | 15.2 | 19.5 | 20.7 | 18.1 | 18.7 | 31.8 | 41.7 |
| mup | 27.0 | 28.3 | 31.7 | 39.9 | 51.9 | 45.1 | 34.9 | 49.2 | 62.3 | 73.0 |
| mwr | 26.4 | 29.7 | 33.9 | 39.2 | 52.2 | 44.6 | 34.6 | 51.3 | 64.2 | 74.5 |
| sat | 8.1 | 6.5 | 8.4 | 9.4 | 27.9 | 16.4 | 17.0 | 16.1 | 20.6 | 45.6 |
| as | 9.5 | 18.1 | 51.1 | 28.5 | 44.1 | 25.9 | 23.6 | 60.0 | 51.7 | 63.3 |
| bho | 23.4 | 25.1 | 24.9 | 33.5 | 42.9 | 40.9 | 32.0 | 51.8 | 55.3 | 61.7 |
| ne | 25.9 | 24.7 | 72.0 | 45.8 | 57.8 | 42.3 | 34.6 | 63.3 | 62.6 | 72.3 |
| or | 9.4 | 5.2 | 45.7 | 19.6 | 52.9 | 19.6 | 17.1 | 63.4 | 56.9 | 68.0 |
| pa | 9.7 | 2.6 | 57.1 | 30.7 | 51.8 | 22.1 | 28.4 | 60.8 | 62.5 | 70.1 |
| ps | 13.3 | 10.9 | 9.2 | 27.3 | 37.6 | 27.7 | 22.0 | 19.4 | 45.7 | 64.5 |
| sa | 16.9 | 18.6 | 15.7 | 29.3 | 34.3 | 31.4 | 25.1 | 32.5 | 51.5 | 59.4 |
| Avg. | 18.1 | 18.6 | 40.8 | 32.1 | 47.5 | 32.7 | 29.2 | 48.4 | 54.5 | 65.1 |
| High Avg. | 18.2 | 17.6 | 67.7 | 36.2 | 56.9 | 31.5 | 33.7 | 59.1 | 59.6 | 68.2 |
| Medium Avg. | 15.4 | 15.0 | 39.4 | 30.7 | 45.9 | 30.0 | 26.1 | 50.2 | 55.2 | 65.6 |
| Low Avg. | 19.5 | 21.3 | 22.9 | 29.9 | 41.9 | 35.0 | 27.7 | 40.0 | 50.5 | 62.6 |
| Lang. Code | XorQA-In-Xx | XorQA-In-En | ||||||||
| LLaMA-65B | Gemma-7B-IT | BLOOMZ-7B | GPT-4 | PaLM-2-L | LLaMA-65B | Gemma-7B-IT | BLOOMZ-7B | GPT-4 | PaLM-2-L | |
| en | 61.1 | 35.5 | 68.7 | 37.4 | 71.4 | 63.9 | 34.7 | 69.7 | 37.9 | 71.5 |
| bn | 15.8 | 9.4 | 0.6 | 16.5 | 40.4 | 16.9 | 25.7 | 68.4 | 46.2 | 55.6 |
| gu | 8.8 | 18.3 | 11.2 | 23.8 | 40.2 | 12.9 | 32.7 | 64.2 | 51.9 | 53.7 |
| hi | 43.5 | 29.6 | 16.5 | 38.3 | 56.5 | 23.1 | 31.1 | 66.1 | 41.6 | 58.8 |
| kn | 8.7 | 19.4 | 8.8 | 28.8 | 41.1 | 9.4 | 26.2 | 63.5 | 54.6 | 59.6 |
| ml | 13.9 | 26.7 | 19.6 | 31.1 | 56.7 | 16.9 | 30.7 | 66.4 | 55.3 | 54.8 |
| mr | 28.1 | 26.4 | 17.7 | 30.4 | 44.5 | 22.3 | 29.1 | 64.0 | 47.9 | 57.4 |
| ta | 16.0 | 17.5 | 10.7 | 25.4 | 40.2 | 19.0 | 26.8 | 64.0 | 50.1 | 57.7 |
| te | 4.0 | 15.3 | 11.4 | 18.5 | 25.6 | 6.0 | 29.3 | 67.6 | 52.0 | 56.3 |
| ur | 20.6 | 7.6 | 1.1 | 18.9 | 31.6 | 21.5 | 34.1 | 58.5 | 44.9 | 61.5 |
| as | 17.6 | 10.8 | 2.0 | 26.4 | 45.5 | 12.5 | 22.3 | 60.0 | 51.6 | 59.3 |
| bho | 19.9 | 9.9 | 2.0 | 23.8 | 32.1 | 21.4 | 26.7 | 55.2 | 46.9 | 59.4 |
| or | 4.0 | 1.6 | 6.2 | 22.1 | 35.9 | 7.7 | 14.8 | 48.0 | 54.6 | 52.6 |
| pa | 6.9 | 13.3 | 4.0 | 25.7 | 40.9 | 10.7 | 30.4 | 63.2 | 47.3 | 55.5 |
| ps | 10.8 | 3.4 | 0.5 | 12.4 | 25.7 | 12.4 | 23.9 | 3.5 | 44.7 | 62.9 |
| sa | 21.9 | 11.0 | 2.1 | 19.0 | 40.3 | 19.0 | 25.5 | 44.8 | 54.8 | 57.8 |
| awa | 26.3 | 13.1 | 7.8 | 28.5 | 41.3 | 23.3 | 25.3 | 58.1 | 50.3 | 58.4 |
| bgc | 18.5 | 16.4 | 10.2 | 20.1 | 28.1 | 23.1 | 25.6 | 55.7 | 43.8 | 58.4 |
| bo | 3.3 | 0.9 | 5.4 | 32.5 | 32.5 | 3.2 | 14.4 | 6.0 | 33.8 | 52.6 |
| brx | 5.3 | 12.7 | 0.9 | 15.5 | 20.0 | 12.1 | 15.6 | 7.6 | 35.9 | 39.1 |
| gbm | 13.7 | 14.9 | 9.5 | 18.5 | 35.4 | 21.6 | 24.9 | 50.7 | 46.1 | 60.2 |
| gom | 15.1 | 8.0 | 0.8 | 19.2 | 33.2 | 14.0 | 22.1 | 37.8 | 41.4 | 54.2 |
| hne | 29.7 | 22.7 | 12.6 | 32.0 | 43.4 | 22.5 | 24.9 | 56.4 | 43.9 | 54.8 |
| hoj | 31.2 | 25.3 | 17.3 | 36.3 | 53.8 | 21.5 | 23.9 | 51.5 | 48.8 | 62.6 |
| mai | 28.0 | 18.9 | 8.0 | 29.4 | 45.1 | 20.5 | 25.9 | 55.2 | 45.0 | 55.1 |
| mni | 8.9 | 4.0 | 0.4 | 13.2 | 23.6 | 9.7 | 14.9 | 17.7 | 36.2 | 47.5 |
| mup | 17.0 | 10.6 | 4.8 | 21.0 | 26.4 | 22.8 | 24.9 | 57.2 | 41.9 | 60.6 |
| mwr | 24.4 | 11.1 | 2.9 | 25.3 | 40.0 | 23.2 | 25.8 | 53.1 | 48.7 | 59.1 |
| sat | 1.3 | 0.3 | 0.0 | 2.6 | 26.6 | 6.7 | 16.7 | 6.2 | 27.5 | 38.6 |
| Avg. | 16.5 | 13.5 | 7.0 | 23.4 | 37.4 | 16.3 | 24.8 | 49.0 | 46.0 | 55.9 |
| High Avg. | 17.7 | 18.9 | 10.8 | 25.8 | 41.9 | 16.4 | 29.5 | 64.7 | 49.4 | 57.3 |
| Medium Avg. | 13.5 | 8.3 | 2.8 | 21.6 | 36.7 | 14.0 | 23.9 | 45.8 | 50.0 | 57.9 |
| Low Avg. | 17.1 | 12.2 | 6.2 | 22.6 | 34.6 | 17.3 | 21.9 | 39.5 | 41.8 | 53.9 |
| XQuAD-In | |||||
| Lang. Code | LLaMA-65B | Gemma-7B-IT | BLOOMZ-7B | GPT-4 | PaLM-2-L |
| en | 62.0 | 45.7 | 86.7 | 64.8 | 83.7 |
| bn | 7.4 | 35.7 | 57.1 | 56.5 | 72.3 |
| gu | 0.5 | 40.6 | 57.0 | 53.9 | 72.1 |
| hi | 25.3 | 49.4 | 63.7 | 63.1 | 76.7 |
| kn | 0.4 | 41.1 | 52.0 | 55.1 | 74.4 |
| ml | 3.6 | 33.7 | 48.8 | 56.5 | 66.6 |
| mr | 14.7 | 33.8 | 58.5 | 57.3 | 76.9 |
| ta | 4.0 | 42.1 | 55.0 | 55.3 | 75.1 |
| te | 0.2 | 36.8 | 51.9 | 48.8 | 68.0 |
| ur | 22.9 | 35.9 | 55.9 | 58.4 | 70.2 |
| as | 3.1 | 28.4 | 48.8 | 53.0 | 66.6 |
| or | 0.7 | 5.9 | 34.0 | 51.1 | 52.0 |
| pa | 1.8 | 40.2 | 61.4 | 59.7 | 60.7 |
| Avg. | 7.1 | 35.3 | 53.7 | 55.7 | 69.3 |
| High Avg. | 8.8 | 38.8 | 55.5 | 56.1 | 72.5 |
| Medium Avg. | 1.9 | 24.8 | 48.1 | 54.6 | 59.8 |
| Lang. Code | XQuAD-In | |
| mT5 XXL | PaLM-2-XS | |
| en | 78.1 | 64.4 |
| bn | 53.2 | 47.7 |
| gu | 53.4 | 21.1 |
| hi | 59.2 | 59.1 |
| kn | 60.2 | 27.9 |
| ml | 52.6 | 30.1 |
| mr | 60.9 | 51.6 |
| ta | 62.8 | 45.1 |
| te | 51.0 | 28.2 |
| ur | 63.2 | 50.9 |
| as | 45.0 | 35.1 |
| or | 25.0 | 6.7 |
| pa | 51.3 | 8.8 |
| Avg. | 53.2 | 34.4 |
| High Avg. | 57.4 | 40.2 |
| Medium Avg. | 40.5 | 16.9 |
| Lang. Code | CrossSum-In | |
| mT5 XXL | PaLM-2-XS | |
| en | 31.8 | 36.6 |
| bn | 25.4 | 28.8 |
| gu | 22.4 | 24.6 |
| hi | 26.2 | 30.5 |
| kn | 26.7 | 28.5 |
| ml | 26.4 | 26.7 |
| mr | 24.4 | 29.8 |
| ta | 31.6 | 32.3 |
| te | 25.8 | 26.9 |
| ur | 26.5 | 28.6 |
| as | 23.9 | 25.8 |
| bho | 22.0 | 24.1 |
| ne | 28.3 | 30.5 |
| or | 23.2 | 24.1 |
| ps | 26.9 | 24.9 |
| pa | 23.4 | 23.7 |
| sa | 25.4 | 26.0 |
| awa | 20.4 | 21.6 |
| brx | 19.7 | 12.6 |
| hne | 22.5 | 23.9 |
| gbm | 18.3 | 18.1 |
| bgc | 22.0 | 23.0 |
| gom | 25.1 | 25.8 |
| mai | 22.1 | 23.8 |
| mup | 24.5 | 25.6 |
| mni | 15.6 | 12.0 |
| mwr | 21.0 | 22.2 |
| hoj | 19.5 | 21.2 |
| sat | 5.9 | 10.0 |
| bo | 7.0 | 5.1 |
| Avg. | 22.5 | 23.5 |
| High Avg. | 26.2 | 28.5 |
| Medium Avg. | 24.7 | 25.6 |
| Low Avg. | 18.7 | 18.8 |
| Lang. Code | XorQA-In-Xx | XorQA-In-En | |||
| mT5 XXL | PaLM-2-XS | mT5 XXL | PaLM-2-XS | ||
| en | 72.4 | 55.8 | 70.7 | 68.1 | |
| bn | 15.4 | 16.3 | 71.9 | 68.6 | |
| gu | 28.7 | 25.9 | 70.6 | 69.0 | |
| hi | 38.3 | 45.8 | 70.6 | 68.5 | |
| kn | 28.2 | 32.0 | 69.3 | 69.3 | |
| ml | 38.2 | 44.0 | 71.1 | 70.5 | |
| mr | 32.8 | 32.9 | 70.5 | 68.9 | |
| ta | 27.2 | 33.8 | 69.0 | 69.8 | |
| te | 21.5 | 24.1 | 70.5 | 70.4 | |
| ur | 28.3 | 18.6 | 68.8 | 66.5 | |
| as | 24.8 | 27.6 | 70.3 | 67.7 | |
| bho | 20.6 | 22.0 | 67.9 | 66.0 | |
| or | 23.9 | 28.6 | 69.1 | 65.3 | |
| pa | 27.3 | 28.5 | 69.0 | 68.5 | |
| ps | 22.4 | 13.1 | 68.3 | 64.2 | |
| sa | 22.1 | 21.5 | 69.1 | 68.0 | |
| awa | 24.2 | 27.7 | 68.0 | 65.2 | |
| bgc | 21.1 | 21.5 | 69.1 | 62.5 | |
| bo | 41.5 | 6.4 | 42.9 | 56.7 | |
| brx | 14.9 | 7.9 | 38.8 | 30.2 | |
| gbm | 15.6 | 18.2 | 65.2 | 62.5 | |
| gom | 17.8 | 21.0 | 64.7 | 64.1 | |
| hne | 27.9 | 32.4 | 68.2 | 64.7 | |
| hoj | 30.8 | 33.3 | 66.2 | 62.1 | |
| mai | 23.5 | 31.8 | 69.2 | 65.0 | |
| mni | 16.8 | 8.9 | 48.6 | 37.4 | |
| mup | 20.1 | 19.1 | 67.1 | 64.2 | |
| mwr | 27.2 | 23.7 | 68.4 | 64.2 | |
| sat | 2.7 | 3.3 | 32.1 | 36.5 | |
| Avg. | 24.5 | 23.9 | 64.8 | 62.7 | |
| High Avg. | 28.8 | 30.4 | 70.3 | 69.1 | |
| Medium Avg. | 23.6 | 23.6 | 68.9 | 66.6 | |
| Low Avg. | 21.9 | 19.6 | 59.1 | 56.6 | |